Please help me solve for "n" in the following equation: 3(5n+7)= -n-11
Thank You

Answer by Boreal(15235)

3(5n+7)= -n-11
distribute
15n + 21 =-n-11
add n to both sides
16n+21= -11
subtract 21 from both sides
16n= -32
n= -2
check
3(-10+7)=-(-2)-11
3(-3)=-9= -(-2)-11